OpenStudy (anonymous):

For each piecewise-defined function find (a) f(-5) (b) f(-1) (c) f(0) (d) f(3) f(x)={x-2 if x<3 5-x if >=3

OpenStudy (anonymous):

a) -5-2=-7 b) -1-2=-3 c) 0-2=-2 d) 5-3=2

OpenStudy (anonymous):

if the x value is less than 3, use the x-2 if the x value is greater than or equal to 3, use 5-x the x value being what you have in the f()